OPTICAL STARK SPECTROSCOPY OF RHODIUM CONTAINING MOLECULES: RhN^{a}.

\underline{TIMOTHY C. STEIMLE}, TONGMEI MA, JAMIE GENGLER, HAILING WANG, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; ZHONG WANG, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973-5000.

Rhodium is very important in both the homogeneous and heterogeneous catalytic induced chemical modification of nitrogen containing molecules. The only practical means of gauging the quantitative predictability of either ab initio and DFT computations used to model catalysis is by a comparison of experimental and predicted properties of bound Rh containging molecules. Here we report on the optical Stark effect in the Q(1) and R(0) lines of the [15.1]1 - X^{1}Σ^{+} (1,0) band of rhodium mononitride, RhN, which were analyzed to determine the permanent electric dipole moments, μ, for the X^{1}Σ^{+} (v=0) and [15.1]1 (v=1) states to be 2.43(5)D and 1.75(1)D, respectively. TJe determined dipole moments are compared with predicted values obtained from DFT^{a} and an all-electron ab initio calculation^{b}. A simple molecular orbital correlation diagram is used to rationalize the relative values of μ for RhN, isovalent IrN^{c} and RhO^{d}. Bonding in the 4d metal containing nitrides, as revealed from an interpretation of μ, will be presented.

aSupported by DoE-Basic Energy Sciences

aStevens et al, Chem. Phys. Lett. 421, 281 (2006).

bShim et al, J. Mol. Struct. 293, 127 (1997).

cA.J. Marr, M.E. Flores and T.C. Steimle, J. Chem. Phys. 104, 8183 (1996).

dJ.Gengler,T. Ma and T.C. Steimle, J. Chem. Phys.(accepted).